OpenStudy (kmullis6):

Help me with my little brother's math problem? I don't think I learned this. :P

OpenStudy (kmullis6):

How can Steve add 48 + 34 by making a ten? What is the sum?

OpenStudy (amistre64):

sounds like common core math

OpenStudy (kmullis6):

I know I feel dumb but I genuinely don't remember learning this way.

OpenStudy (amistre64):

add 2 add 6, then subtract 2 and 6

OpenStudy (amistre64):

not too sure about the method myself

rishavraj (rishavraj):

i guess it would be 40 + 30 = 70 8 + 4 = 12 and finally add 70 + 12

OpenStudy (amistre64):

48 + 34 (40+8) + (30+4) (40+30} + (8+4) (40+30} + (8+2) +2

OpenStudy (kmullis6):

This is the example: Add 26 + 17 by breaking apart numbers to make a ten. Use a number that adds with the 6 in 26 to make a 10. Since 6 + 4 = 10, use 4. Think: 17 = 4 + 13 Add 26 + 4 = 30 Add 30 + 13 = 43 So, 26 + 17 = 43

OpenStudy (kmullis6):

I don't know if it's just late and I can't wrap my head around what the ten's use is?

OpenStudy (kmullis6):

@amistre64

OpenStudy (amistre64):

the 10s method is spose to be based on the assumption that counting by 10s is simpler to accomplish so we deconstruct the problem into counts of 10, then add in whats left

OpenStudy (amistre64):

or someting along those lines

OpenStudy (nikato):

i dont get why anyone would use this method but... 48+34 think: 48 is closer to 50 and you know that 48+2=50 and 2+32=34 so just add 50+32 to get your anwer

OpenStudy (anonymous):

I don't even know why they still teach common core....

OpenStudy (amistre64):

its all about the benjamins ...

OpenStudy (kmullis6):

Haha ok so I'm not the only one :P

OpenStudy (anonymous):

No you arnt, Kmullis.